The following selected information was extracted from the 20x1 accounting records of Lone Oak Products:

Raw material purchases $ 176,000
Direct labor 256,000
Indirect labor 108,000
Selling and administrative salaries 133,000
Building depreciation* 80,000
Other selling and administrative expenses 195,000
Other factory costs 343,000
Sales revenue ($130 per unit) 1,495,000

*Seventy-five percent of the company's building was devoted to production activities; the remaining 25 percent was used for selling and administrative functions.


Inventory data:

January 1 December 31
Raw material $

15,900

$

18,200

Work in process

35,800

62,100
Finished goods*

113,100

97,900

*The January 1 and December 31 finished-goods inventory consisted of 1,350 units and 1,190 units, respectively.

a. Calculate Lone Oak’s manufacturing overhead for the year.

b. Calculate Lone Oak’s cost of goods manufactured.

c. Compute the company’s cost of goods sold.     

d. Determine net income for 20x1, assuming a 30% income tax rate.

e. Determine the number of completed units manufactured during the year.
Answer #1

a.

Calculation of Lone Oak's Manufacturing Overhead for the year:

Indirect labor

$      108,000

Building Depreciation = 80000*75% =

$        60,000

Other Factory Costs

$      346,000

Lone Oak's Manufacturing Overhead for the year =

$ 511,000

b.

Calculation of Lone Oak's Cost of Goods Manufactured for the year:

Direct Material :

Raw Material Purchase for the year

$      176,000

Add: Raw Material Inventory Jan. 1, 20x1

$        15,900

Less: Raw Material Inventory Dec. 31, 20x1

$      (18,200)

Total Direct Material used

$ 173,700

Direct Labor

$ 256,000

Manufacturing Overhead

$ 511,000

Total Manufacturing Costs

$ 940,700

Add: Work in Process inventory, Jan 1, 20x1

$    35,800

Less: Work in Process inventory, Dec 31, 20x1

$ (62,100)

Total Cost of Goods Manufactured

$ 914,400

c.

Calculation of Lone Oak's Cost of Goods sold for the year:

Total Cost of Goods Manufactured

$ 914,400

Add: Finished Goods inventory, Jan 1, 20x1

$ 113,100

Less: Finished Goods inventory, Dec 31, 20x1

$ (97,900)

Lone Oak's Cost of Goods sold for the year =

$ 929,600

d

Calculation of Net income:

Sales Revenue

$ 1,495,000

Less: Cost of Goods sold

$   (929,600)

Gross Revenue

$ 565,400

Less: Selling and administrative Expenses:

Selling and administrative Salaries

$      133,000

Building Depreciation = 80000*25% =

$        20,000

Other Selling and administrative expenses

$      195,000

Total Selling and administrative Expenses

$ 348,000

Income before tax

$ 217,400

Less: Tax @ 30%

$ (65220)

Net Income

$ 152,180

e. Number of units manufactured during the year:
Sales [1,495,000/130] 11500
Add: ending inventory 1190
Less: beginning inventory (1350)
Number of units manufactured 11340